i need detailed solutions with step by step explanations.If a regression model of the formy = β0+Pix! + +p6x6 is fit to 109 observations on each variable and yields an Rvalue of 0.4
Answer #1

Nk that S.CA SSK = Σ($-37 Sivce aurl SS T SST n- SST 27 0:48 SS T =SS T-306 ラ SS T = 588-462 588.462- 588·46 2-306-28 2·46 2.C@ose6 = 15.692 rok Mar note at MSR- SSR 6 dF(SSR) = SSE df(ssE) to and F=-MSK MSE Arova table-u- as llows:- F-stat atuat 41